# Experimental observation of gapped shear waves and liquid-like to gas-like dynamical crossover in active granular matter

### v1

@@ Abstract
Unlike crystalline solids, liquids do not exhibit long-range order. Their atoms undergo frequent structural rearrangements, resulting in the long-wavelength dynamics of shear fluctuations in liquids being diffusive, rather than propagating waves, as observed in crystals. When considering shorter time and length scales, molecular dynamics simulations and theoretical propositions suggest that collective shear excitations in liquids display a gap in wave-vector space, referred to as the $k$-gap. Above this gap, solid-like transverse waves re-emerge. However, direct experimental verification of this phenomenon in classical liquids remains elusive, with the only documented evidence from studies in two-dimensional dusty plasmas. Active granular systems provide a novel platform for exploring the emergence of collective dynamics and showcasing a rich interplay of complex phases and phenomena. Our study focuses on bi-disperse active Brownian vibrators. Through measurements of the pair correlation functions, mean square displacements, velocity auto-correlation functions, vibrational density of states, and a detailed analysis of microscopic atomic motion, we demonstrate that this active system exhibits both gas-like and liquid-like phases, depending on the packing fraction, despite pure hard-disk-like repulsive interactions. Notably, within the granular liquid-like phase, we experimentally validate the existence of a $k$-gap in the dispersion of transverse excitations. This gap becomes more significant with a decrease in packing fraction and disappears into the gas phase, aligning with theoretical expectations. Our results offer a direct experimental confirmation of the $k$-gap phenomenon, extending its relevance beyond classical thermal liquids to active granular systems, and reveal the existence of intriguing similarities between the physics of active granular matter and supercritical fluids.
@@ Introduction
Because of the random atomic distribution and the absence of a fixed equilibrium reference frame, the fate of phonons and the vibrational properties of liquids represent a much harder challenge for both theory and experiments [3]. In liquids, the dynamics of longitudinal long-wavelength fluctuations have been experimentally ascertained [4] to be qualitatively identical to that of solids, despite a smaller sound speed. On the contrary, the dynamics of transverse (or shear) long-wavelength fluctuations are radically different. As an immediate consequence of the absence of long-range order, or equivalently the absence of spontaneous translational symmetry breaking, liquids have a vanishing static shear modulus and, at small wave-vector, they display a shear diffusion mode rather than propagating shear waves (transverse phonons) as in solids [1].
On the basis of Maxwell and Frenkel original intuition [5], shear fluctuations in liquids have been predicted to obey the following telegrapher equation,
$\omega_{T}^{2}+i\omega_{T}\tau^{-1}=v_{T}^{2}k^{2}.$ (1)
Similar conclusions have been later obtained using generalized hydrodynamics [6, 7, 8], in which higher-$k$ corrections can also be systematically included. In (1), $v_{T}$ is the transverse speed of sound related to the instantaneous shear modulus $G_{\infty}$ and $\tau$ a characteristic timescale identified with the Maxwell relaxation time $\tau_{m}=\eta/G_{\infty}$ (where $\eta$ is the shear viscosity). By solving (1) (see [9] for an extensive review), the dispersion of shear waves is obtained,
@@
For long-wavelengths, one recovers the hydrodynamic shear diffusion mode with collective diffusion constant $D_{c}\equiv v^{2}_{T}\tau=\eta/\rho$ (with $\rho$ the mass density of the system) predicted by Navier-Stokes equations [10]. Importantly, above a critical wave-vector $k_{g}$, known as $k$-gap, the real part of the frequency becomes nonzero. Above $k_{g}$, when the real part of $\omega_{T}$ becomes larger than its imaginary part, propagating solid-like shear waves are then expected to emerge in liquids. This leads to a corresponding elastic-like response below a certain critical distance, $L_{c}\equiv 2\pi/k_{g}$. In simpler terms, it means that liquids are predicted to exhibit solid-like transverse vibrational modes not only for high-frequency, $\omega\gg 1/\tau$ (as proposed initially by Frenkel [5]), but also for large wave-vectors $k\gg k_{g}$.
The $k$-gap is expected to appear at the melting temperature and to expand into the liquid phase as the temperature increases [11]. It then reaches the maximum wave vector allowed at the edge with the gas phase. The existence of the $k$-gap, and its properties as described by equation (1), have been confirmed by several molecular simulations of classical liquids [11, 12, 13, 14, 15, 16, 17] and other liquid systems [18, 19, 20], indicating the validity of the theoretical framework, even from a quantitative perspective [21]. Additionally, the similarity between vibrational modes at high frequency/wave-vector in liquids and solids [22, 23, 24], and the solid-like nature of confined liquids at low frequencies, as predicted by the concept of an elastic critical length $L_{c}$, have been experimentally verified [25, 26]. However, due to the limitations of experimental scattering techniques at low $k$ and $\omega$, the $k$-gap has not been observed experimentally in classical liquids. The only recorded observation of the $k$-gap has been in complex dusty plasmas [27], which are systems of charged particles that interact strongly via Coulomb forces and exhibit liquid-like collective dynamics [28].
@@
The densely packed configurations of granular materials often exhibit fluid-like behaviors when subjected to external forces [30, 31, 32, 33, 34, 35]. However, due to the presence of permanent contacts and force chains [36], measuring the Hessian matrix directly in densely packed granular matter has proven to be a challenging task, even in the quasi-static limit [37, 38, 39, 40]. In contrast, loose granular matter, where collisions primarily govern particle interactions, has seen theoretical analyses of hydrodynamics and collective modes in granular fluids. This involves formulating transport equations for essential hydrodynamic quantities like mass, momentum, and heat, followed by a linear stability analysis of the homogeneous states [41, 42, 43, 44, 45, 46, 47]. While transverse modes decouple [45], the longitudinal sector becomes intricate due to the non-conservation of energy, leading to significant modifications in the longitudinal channel. Notably, to the best of our knowledge, the discussion of the $k$-gap in the hydrodynamics of granular fluids has been absent, as it extends beyond the conventional long-wavelength hydrodynamic description.
@@
If the k-gap description can apply to granular fluids, then a homogeneously driven granular fluid would be the most straightforward scenario to explore. However, creating a homogeneously driven experimental system has proven challenging [48, 49, 50] due to the influence of gravity and anisotropic driving in three-dimensional (3D) systems or the implementation of boundary driving in quasi-two-dimensional (2D) vertical systems [51, 52]. Additionally, the influence of 3D effects in quasi-2D horizontal systems has made it difficult to achieve homogeneous driving [53, 54, 55]. While a few quasi-2D systems have achieved homogeneous driving [56, 57, 48], some lacked single-particle velocity Gaussian statistics [56, 57]. In contrast, others incorporated persistent unidirectional rotation at the single-particle level [48], introducing additional complexities in energy injection at the single-particle scale. Recently, Chen et al. designed an experimental system that achieves homogeneous driving, single-particle velocity and rotation statistics with Gaussian distributions of zero means in a quasi-2D system [49, 50]. This system closely aligns with the active Brownian particles introduced in theoretical studies from the perspective of active matter [58].
In this work, we conducted an experimental study on the structure, single-particle scale dynamics, and collective dynamics of active granular matter. We controlled the packing fraction and explored the entire phase diagram. Our findings indicate the presence of a liquid-like to gas-like dynamical crossover, similar to that observed in supercritical fluids [59]. This crossover is well-correlated with the experimentally measured vibrational density of states. Additionally, we observed the dispersion of gapped-shear waves in the liquid phase, where the $k$-gap increases as the packing fraction decreases.
@@
Fig.1 displays the top view of a layer of bi-disperse Brownian vibrators positioned on the surface of a shaker. This layer of particles is confined within a flower-shaped boundary to prevent the creep motion of particles near the boundary. Each Brownian vibrator has a flat, disk-shaped cap with twelve alternatively inclined legs below the cap. When a vertical sinusoidal vibration is applied to the supporting base, each single Brownian vibrator performs 2D Brownian motion. Previous studies by Chen et al. revealed that the translational and rotational velocities of a single Brownian vibrator follow Gaussian distributions with zero means [49]. These features lead us to term the particle an “active Brownian particle”, closely mimicking the conditions studied in theoretical investigations of active matter systems [58]. Moreover, in a collection of Brownian vibrators of the same size, the translational velocities of individual particles follow a Maxwell-Boltzmann distribution for low and intermediate speeds, but show high-energy tails that deviate significantly from the Maxwell-Boltzmann distribution for large speeds, which can be attributed to the inelastic collisions of particles and the homogeneous driving [49]. Unlike the mono-disperse systems studied earlier [49, 50], the present system is bi-disperse, which prevents crystallization at high packing fractions, as depicted in Fig.1. More details about this setup are presented in the Methods.
@@
The VACF of a gas decreases continuously with time, while for liquids near melting point and solids, it shows a combination of an oscillatory and a decaying term. The presence of an oscillatory part in the $C(t)$ can be identified by looking for the occurrence of a minimum or a change in its slope. For high packing fractions, $C(t)$ displays a clear minimum below $t\approx 0.1s$ that gradually disappears as the packing fraction decreases, as shown in Fig.2(b). At a packing fraction $\phi=0.618$, the minimum in the VACF is no longer present, and $C(t)$ becomes a continuously decreasing function, as expected in a gas. In the field of supercritical fluids, this dynamical transition determines the so-called Frenkel line that separates the rigid liquid phase, presenting oscillatory motion, and the non-rigid gas-like fluid phase. Evidence for the structural nature of the Frenkel line, hinting towards a possible equivalence with the Fisher-Widom line concept, has been reported in supercritical fluids [59]. Despite the complete equivalence between the structural and dynamical criteria remains unproved, a direct connection between the dynamical crossover and thermodynamics has been demonstrated [62]. Aware of these distinctions, in the rest of the manuscript, we will adopt the jargon rigid (liquid-like) and non-rigid (gas-like) states interchangeably. According to Frenkel’s theory [5], this dynamical crossover corresponds microscopically to a situation where the jumping time between oscillatory motion around different local minima of the potential becomes comparable with the shortest vibration time. This dynamical crossover is also expected to coincide with the disappearance of collective shear waves at all frequencies in the liquid. Notably, the critical packing fraction of $\phi=0.618$ is very close to the value at which medium-range order disappears in the pair correlation functions shown in panel (a). This suggests a significant link between structure and collective dynamics in active granular systems, similar in spirit to the results presented in [59].
@@
The MSD and the VACF are closely related to each other,
@@
where in the second line, the factor of four arises because of a combination of time-reversal symmetry $t\rightarrow-t$ and the dimensionality of the system, and the last line is obtained using the change of variable from $t^{\prime\prime}$ to $s=t^{\prime}-t^{\prime\prime}$ [63].
The above Eq. (4) yields the direct connection between the second time derivative of $MSD$ and $Z(t)$ as $\frac{d^{2}}{dt^{2}}MSD(t)=4Z(t)$, which proves the equivalence of MSD and VACF [59]. Upon comparing panels b and d of Fig.2, we experimentally verified this equivalence, and the results show an excellent quantitative mutual agreement.
Figure 3: The role of the packing fraction in active granular matter. The average velocity squared $Z(0)$ as a function of the packing fraction $\phi$. The empty circles indicate the experimental data. The solid blue line is the phenomenological function $Z(0)=9.73\sqrt{0.8318-\phi}$.
Both structural and dynamical observables indicate that the packing fraction $\phi$ of the Brownian vibrators plays the role of temperature in thermal fluids, driving the system from a gas-like to a liquid-like phase. In order to provide further qualitative evidence for this analogy, in Fig.3, we plot the average velocity squared $Z(0)$, that is proportional to the average kinetic energy, as a function of the packing fraction. We observe a clear anti-correlation between the two quantities, which is well fitted by a phenomenological function $Z(0)\propto\sqrt{\phi_{c}-\phi}$, with $\phi_{c}\approx 0.8318$. In the literature (e.g., [64]), $Z(0)$ has been often associated to an effective granular temperature, $T_{e}$. Our results demonstrate therefore that $T_{e}$ anti-correlates with $\phi$, consistent with the dynamics experimentally observed both at the particle-level and collective scale. It is interesting to notice that the scaling of the experimental data follows a mean-field behavior. We are not aware of any theoretical explanation of this phenomenon. Finally, this analysis suggests the existence of a critical packing fraction $\phi_{c}$ which might be connected to a jamming type transition in active granular systems. We leave the exploration of these two interesting points for future research.
@@
In Fig.4, panel (a), we observe that for the lowest packing fraction data, the VDOS decreases monotonically with frequency and can be accurately described by a Lorentzian line shape, $g(\omega)=\alpha/(\alpha^{2}+\omega^{2})$ (dashed black line), at least for frequencies below 5 Hz. This line shape is indicative of purely Langevin diffusive dynamics. It is typical of a gas-like state, where atomic collisions are almost uncorrelated and independent and can be described by kinetic theory. The Lorentzian fit becomes less accurate at higher frequencies, indicating that the low-packing fraction system is a dilute liquid rather than an ideal gas of free particles. Additionally, the longitudinal and transverse components of $g(\omega)$ are identical, confirming the emergent isotropy of the low-packing fraction phase. Due to the dilute packing, the constituent particles exhibit random, uncorrelated motion in both amplitude and direction. Individual particles’ motion is collisional, resulting in substantial displacements away from their initial positions.
[…]
@@
Moving further to $\phi=0.811$ (panel (c) in Fig.4) results in the disappearance of the Lorentzian gas-like contribution to the VDOS at low frequency. The total VDOS increases monotonically with frequency up to a peak located around $\omega=15$ Hz, which corresponds approximately to the average pseudo-Van Hove energy of the emergent collective longitudinal and transverse excitations, as shown in Fig.5. Notably, the zero-frequency values of the total VDOS $g(0)$ exhibit a substantial decrease with an increase in the packing fraction $\phi$. This value correlates with the self-diffusion constant $D$, and a direct relation between $D$ and $g(0)$ can be derived for pure Langevin diffusion. This is consistent with the experimental data for the MSD presented in panel (c) of Fig.2. It is also important to note that $g(0)$ is predominantly dominated by the transverse component, and $g_{L}(0)$ is almost zero at $\phi=0.811$. Additionally, the system becomes strongly anisotropic as the longitudinal and transverse VDOS are significantly different, and higher-energy modes appear in the spectrum up to a frequency of approximately $35$ Hz.
@@
As the packing fraction of the system increases to $\phi=0.822$, the value of $g(0)$ becomes extremely small but still finite, indicating that the system is close to the solid phase but not quite there yet. This is later confirmed by the dispersion of shear waves. The longitudinal and transverse VDOS are linear in frequency up to about 8 Hz, as dictated by Debye’s law in 2D, and also commonly found in bulk liquids [66, 67]. However, a sharper peak appears around 12 Hz, which is attributed to the flattening of the dispersion of the collective modes obtained from the dynamical structure factor (panels (d) and (h) in Fig.5). As a result, atomic displacements become small, and granular particles move very little away from their initial positions. The dynamics and corresponding VDOS increasingly resemble those of a dense viscous liquid with enhanced solid-like elastic vibrations or a liquid near melting.

### v2

# +]These authors contributed equally to this work

@@ Abstract
Unlike crystalline solids, liquids lack long-range order, resulting in diffusive shear fluctuations rather than propagating waves. Simulations predict that liquids exhibit a $k$-gap in wave-vector space, where solid-like transverse waves reappear above this gap. Experimental evidence in classical liquids has been limited, observed only in 2D dusty plasmas. Here, we investigate this phenomenon using active Brownian vibrators and uncover distinct gas-like and liquid-like phases depending on the packing fraction. We measure key properties, including pair correlation functions, mean square displacements, velocity auto-correlation functions, and vibrational density of states. In the liquid-like phase, we confirm the $k$-gap in transverse excitations, whose size grows as the packing fraction decreases and eventually disappears in the gas phase. Our findings extend the concept of the $k$-gap to active granular systems and reveal striking parallels with supercritical fluids.
keywords:
active granular matter, gapped shear wave
@@ Introduction
Because of the random atomic distribution and the absence of a fixed equilibrium reference frame, the fate of phonons and the vibrational properties of liquids represent a much harder challenge for both theory and experiments [3]. In liquids, the dynamics of longitudinal long-wavelength fluctuations have been experimentally ascertained [4] to be qualitatively identical to that of solids, despite a smaller sound speed. On the contrary, the dynamics of transverse (or shear) long-wavelength fluctuations are radically different. Liquids have a vanishing static shear modulus and, at small wave-vector, they display a shear diffusion mode rather than propagating shear waves (transverse phonons) as in solids [1].
Leveraging on a simple viscoelastic model, Maxwell [5] proposed that shear stress in liquids has a characteristic exponential decay time $\tau_{M}\equiv\eta/G_{\infty}$, where $\eta$ is the shear viscosity and the instantaneous shear modulus $G_{\infty}$. This timescale is now known as the Maxwell relaxation time. Based on a more microscopic picture of liquid dynamics, Frenkel later proposed [6] to identify such a timescale with the time of local particle re-arrangements, corresponding to hopping processes over potential barriers.
The emerging Maxwell-Frenkel picture of liquid dynamics (see [7] for the complete history lesson) suggest that shear fluctuations in liquids obey the following telegrapher equation,
$\omega_{T}^{2}+i\omega_{T}\tau_{M}^{-1}=v_{T}^{2}k^{2}.$ (1)
In Eq. (1), $v_{T}$ is the transverse speed of sound related to the instantaneous shear modulus $G_{\infty}$. By solving (1) (see [8] for an extensive review), the dispersion of shear waves is obtained,
@@
For long-wavelengths, one recovers the hydrodynamic shear diffusion mode with collective diffusion constant $D_{c}\equiv v^{2}_{T}\tau=\eta/\rho$ (with $\rho$ the mass density of the system) predicted by Navier-Stokes equations [9]. Above a critical wave-vector $k_{g}$, known as $k$-gap, the real part of the frequency becomes nonzero. Above $k_{g}$, when the real part of $\omega_{T}$ becomes larger than its imaginary part, propagating solid-like shear waves are then expected to emerge in liquids. This leads to a corresponding elastic-like response below a certain critical distance, $L_{c}\equiv 2\pi/k_{g}$. In simpler terms, it means that liquids are predicted to exhibit solid-like transverse vibrational modes not only for high-frequency, $\omega\gg 1/\tau$ (as proposed initially by Frenkel [6] based on a single particle picture), but also for large wave-vectors $k\gg k_{g}$.
The $k$-gap is expected to appear at the melting temperature and to expand into the liquid phase as the temperature increases [10]. It then reaches the maximum wave vector allowed at the edge with the gas phase. The existence of the $k$-gap, and its properties as described by equation (1), have been confirmed by several molecular simulations of classical liquids [10, 11, 12, 13, 14, 15, 16, 17] and other liquid systems [18, 19, 20], indicating the validity of the theoretical framework, even from a quantitative perspective [21]. Additionally, the similarity between vibrational modes at high frequency/wave-vector in liquids and solids [22, 23, 24], and the solid-like nature of confined liquids at low frequencies, as predicted by the concept of an elastic critical length $L_{c}$, have been experimentally verified [25, 26]. However, due to the limitations of experimental scattering techniques at low $k$ and $\omega$, the $k$-gap has not been observed experimentally in classical liquids. The only recorded observation of the $k$-gap has been in complex dusty plasmas [27], which are systems of charged particles that interact strongly via Coulomb forces and exhibit liquid-like collective dynamics [28].
@@
The densely packed configurations of granular materials often exhibit fluid-like behaviors when subjected to external forces [30, 31, 32, 33, 34, 35]. However, due to the presence of permanent contacts and force chains [36], measuring the Hessian matrix directly in densely packed granular matter has proven to be a challenging task, even in the quasi-static limit [37, 38, 39, 40]. In contrast, loose granular matter, where collisions primarily govern particle interactions, has seen theoretical analyses of hydrodynamics and collective modes in granular fluids. This involves formulating transport equations for essential hydrodynamic quantities like mass, momentum, and heat, followed by a linear stability analysis of the homogeneous states [41, 42, 43, 44, 45, 46, 47]. While transverse modes decouple [45], the longitudinal sector becomes intricate due to the non-conservation of energy, leading to significant modifications in the longitudinal channel. To the best of our knowledge, the discussion of the $k$-gap in the hydrodynamics of granular fluids has been absent, as it extends beyond the conventional long-wavelength hydrodynamic description. From an experimental standpoint, the primary advantage of granular fluids over traditional molecular liquids is the macroscopic size of the particles, which are measured in centimeters in this case. This larger size significantly facilitates the tracking of particle positions and dynamics using cameras. In contrast, such tracking is extremely challenging, if not impossible, with molecular liquids. In those cases, excitations can only be investigated through techniques like X-ray or inelastic neutron scattering, which are far more complicated, especially in the frequency and wave-vector ranges where the $k$-gap is expected to emerge.
@@
If the k-gap description can apply to granular fluids, then a homogeneously driven granular fluid would be the most straightforward scenario to explore. However, creating a homogeneously driven experimental system has proven challenging [48, 49, 50] due to the influence of gravity and anisotropic driving in three-dimensional (3D) systems or the implementation of boundary driving in quasi-two-dimensional (2D) vertical systems [51, 52]. Additionally, the influence of 3D effects in quasi-2D horizontal systems has made it difficult to achieve homogeneous driving [53, 54, 55]. While a few quasi-2D systems have achieved homogeneous driving [56, 57, 48], some lacked single-particle velocity Gaussian statistics [56, 57]. In contrast, others incorporated persistent unidirectional rotation at the single-particle level [48], introducing additional complexities in energy injection at the single-particle scale. Recently, Chen et al. designed an experimental system that achieves homogeneous driving, single-particle velocity and rotation statistics with Gaussian distributions of zero means in a quasi-2D system [49, 50]. This system closely aligns with the active Brownian particles introduced in theoretical studies from the perspective of active matter [58]. We notice that our experimental setup can be considered as an active system since the external energy input, that maintains the system out of equilibrium, acts individually and independently on each “active particle” [59].
Compared to a dusty plasma, a nonequilibrium system made up of micron-sized charged particles suspended in a plasma, an active granular system differs significantly in its interaction potential and driving mechanism. In dusty plasmas, the potential is governed by a Yukawa potential, while in active granular systems, interactions occur through inelastic collisions and solid friction. Additionally, the driving mechanisms differ: dusty plasma is driven by laser heating at the boundaries, while active granular systems experience homogeneous and random driving of individual particles. Our experimental findings add to the previous experimental observation of a $k$-gap dispersion in a dusty plasma [27], demonstrating the universality of the $k$-gap phenomenon in the liquid phases of matter.
Active granular systems provide a novel platform for exploring the emergence of collective dynamics and showcasing a rich interplay of complex phases and phenomena. Our study focuses on bi-disperse active Brownian vibrators. Through measurements of the pair correlation functions, mean square displacements, velocity auto-correlation functions, vibrational density of states, and a detailed analysis of particle motion, we demonstrate that this active system exhibits both gas-like and liquid-like phases, depending on the packing fraction, despite pure hard-disk-like repulsive interactions. Within the granular liquid-like phase, we experimentally validate the existence of a $k$-gap in the dispersion of transverse excitations. This gap becomes more significant with a decrease in packing fraction and becomes ill-defined in the gas phase because of the disappearance of well-defined modes, aligning with theoretical expectations. Our results offer a direct experimental confirmation of the $k$-gap phenomenon, extending its relevance beyond classical thermal liquids to active granular systems, and reveal the existence of similarities between the physics of active granular matter and supercritical fluids.
Result and discussions
@@
Fig.1 displays the top view of a layer of bi-disperse Brownian vibrators positioned on the surface of a shaker. This layer of particles is confined within a flower-shaped boundary to prevent the creep motion of particles near the boundary. Each Brownian vibrator has a flat, disk-shaped cap with twelve alternatively inclined legs below the cap. When a vertical sinusoidal vibration is applied to the supporting base, each single Brownian vibrator performs 2D Brownian motion. Previous studies by Chen et al. revealed that the translational and rotational velocities of a single Brownian vibrator follow Gaussian distributions with zero means [49]. These features lead us to term the particle an “active Brownian particle”, closely mimicking the conditions studied in theoretical investigations of active matter systems [58]. Moreover, in a collection of Brownian vibrators of the same size, the translational velocities of individual particles follow a Maxwell-Boltzmann distribution for low and intermediate speeds, but show high-energy tails that deviate significantly from the Maxwell-Boltzmann distribution for large speeds, which can be attributed to the inelastic collisions of particles and the homogeneous driving [49]. Unlike the mono-disperse systems studied earlier [49, 50], the present system is bi-disperse, which prevents crystallization at high packing fractions, as depicted in Fig.1. More details about this setup are presented in the Methods. We emphasize that the frequency of the vertical sinusoidal vibration applied is of $100$ Hz. In the rest of the manuscript, we will consider only time-scales which are parameterically longer than the driving frequency. In that regime, equilibrium thermodynamic and hydrodynamic concepts can still be applied.
Unlike conventional polar particles studied previously [60, 61, 62], the lack of particle-scale built-in asymmetry makes our experimental system unique and novel. The active force on each nonpolar particle results from collisions between the tilted legs and the vibrating bottom surface, causing the particle’s central axis to tilt slightly away from gravity, allowing only some legs—often represented as a single leg—to be propelled upon contact with the surface. These interactions produce minimal correlation in the contact angle, leading to random driving force directions while maintaining a nearly constant magnitude (see Refs. [49, 50] for details). In this system, there is no fluid flow from a surrounding solvent, and the main dissipation with the environment results from friction and inelastic collisions. Consequently, our system naturally belongs to the category of dry active matter.
In our present system of bi-disperse (nonpolar) Brownian vibrators, we have observed no global flocking within the several-hour experimental time window, which is likely due to the disorder introduced by bidispersity, in contrast to our previous monodisperse systems [50], where we observed global flocking, aligning with the theoretical investigation of active Brownian particles [58]. Furthermore, we have not observed any phase separation, unlike the self-propelled binary colloids known as Quincke rollers, where significant demixing of small and large colloidal particles occurs after the system begins a global rotational collective motion [63]. On a microscopic level, the interactions between these Quincke rollers are influenced by electrostatic and hydrodynamical forces, which are quite different from the interactions observed in our nonpolar granular disks in a dry environment. While examining phase separation in nonpolar disks is an promising area for future research, it is currently beyond the scope of this study. Real-time videos of the particle motion are provided with from the Supplementary Movie 1 to 4 for different packing fraction to confirm these statements.
@@
The VACF of a gas decreases continuously with time, while for liquids near melting point and solids, it shows a combination of an oscillatory and a decaying term. The presence of an oscillatory part in the $C(t)$ can be identified by looking for the occurrence of a minimum or a change in its slope. For high packing fractions, $C(t)$ displays a clear minimum below $t\approx 0.1s$ that gradually disappears as the packing fraction decreases, as shown in Fig.2(b). At a packing fraction $\phi=0.618$, the minimum in the VACF is no longer present, and $C(t)$ becomes a continuously decreasing function, as expected in a gas. In the field of supercritical fluids, this dynamical transition determines the so-called Frenkel line that separates the rigid liquid phase, presenting oscillatory motion, and the non-rigid gas-like fluid phase. Evidence for the structural nature of the Frenkel line, hinting towards a possible equivalence with the Fisher-Widom line concept, has been reported in supercritical fluids [66]. Despite the complete equivalence between the structural and dynamical criteria remains unproved, a direct connection between the dynamical crossover and thermodynamics has been demonstrated [67]. Aware of these distinctions, in the rest of the manuscript, we will adopt the jargon rigid (liquid-like) and non-rigid (gas-like) states interchangeably. According to Frenkel’s theory [6], this dynamical crossover corresponds microscopically to a situation where the jumping time between oscillatory motion around different local minima of the potential becomes comparable with the shortest vibration time. This dynamical crossover is also expected to coincide with the disappearance of collective shear waves at all frequencies in the liquid. Frenkel idea relies on a single particle picture, while the $k$-gap equation (1) describes collective dynamics. The time-scale $\tau$ appearing in Eq. (1) is therefore more correctly identified with the Maxwell relaxation time. In simple fluids the Maxwell time is very close to the lifetime of local connectivity [68], that is another single particle concept that will be analyzed below. The critical packing fraction of $\phi=0.618$ is very close to the value at which medium-range order disappears in the pair correlation functions shown in panel (a). This suggests a significant link between structure and collective dynamics in active granular systems, similar in spirit to the results presented in [66]. We notice that there is now firm experimental evidence that the change of particle dynamics at the Frenkel line is seen in structural changes [69]. Despite a complete analysis of these structural changes being beyond the scope of the present work, in the Supplementary note 2, we provide a preliminary study of the experimental structure factor $S(k)$ for four different packing fractions that confirms the gradual disappearance of structural order by decreasing the packing fraction. The relation between structural changes and dynamics in liquids is still poorly understood (see, for example, [7]). It would be necessary to explore this connection further in granular fluids.
@@
The MSD and the VACF are closely related to each other since $\frac{d^{2}}{dt^{2}}\mathrm{MSD}(t)=4Z(t)$ [70]. This equivalence has been already utilized in Ref. [66] to investigate the “liquid-gas” transition. Upon comparing panels b and d of Fig.2, we experimentally verified this equivalence, and the results show an excellent quantitative mutual agreement.
Figure 3: The role of the packing fraction in active granular matter. The average velocity squared $Z(0)$ for small (red), large (blue) and all (gray) particles as a function of the packing fraction $\phi$. All the data are well fitted by a phenomenological function $Z(0)=a\sqrt{b-\phi}$, where $b=0.8318$ for all curves. The error bar of averaged value here is obtained as the difference of minimum and averaged value.
Granular systems are athermal in nature because of the macroscopic size of their constituents. At fixed activity, it is therefore necessary to understand which parameter plays an analogous role of temperature in thermal fluids, driving the system from a gas-like to a liquid-like phase. Both structural and dynamical observables suggest that the packing fraction $\phi$ of the Brownian vibrators plays such a role. In order to provide further qualitative evidence for this analogy, in Fig.3, we plot the average velocity squared $Z(0)$, that is proportional to the average kinetic energy, as a function of the packing fraction for small (red), large (blue) and all (gray) particles. We observe a clear anti-correlation between the two quantities, which is well fitted by a phenomenological function $Z(0)\propto\sqrt{\phi_{c}-\phi}$, with $\phi_{c}\approx 0.8318$. In the literature (e.g., [71]), $Z(0)$ has been often associated to an effective granular temperature, $T_{e}$. Our results demonstrate therefore that $T_{e}$ anti-correlates with $\phi$, consistent with the dynamics experimentally observed both at the particle-level and collective scale. Temperature is clearly defined only in equilibrium thermodynamic systems. Defining temperature rigorously in non-equilibrium systems is often very challenging. Here, we introduce the term ‘temperature’ for our system only in a vague, intuitive sense, drawing an analogy with well-defined meanings for ordinary materials made of molecules in thermal equilibrium. It is necessary to notice that the scaling of the experimental data follows a mean-field behavior. We are not aware of any theoretical explanation of this phenomenon. This analysis suggests the existence of a critical packing fraction $\phi_{c}$ which might be connected to a jamming type transition in active granular systems. We leave the exploration of these two points for future research. Finally, we notice that the average velocity squared $Z(0)$ for large particles is consistently larger than that of small particles, with this difference becoming more pronounced for small packing fraction. Given that the steady state velocity is a result of the balance between the energy injected and the energy loss caused by friction, this might be explained by the fact that larger particles experience stronger drag force induced by activity.
@@
The radial pair correlation function is defined as,
$g(r)=\dfrac{1}{2\pi rN\rho}\sum_{i=1}^{N}\sum_{j\neq i}^{N}\delta\left(r-|\vec{r}_{ij}|\right),$ (4)
where $\vec{r}_{ij}$ is the vector between $i$th and $j$th particles, $N$ is the total number of particles, and $\rho$ is the particle density of system. As shown in Fig.2 a, three peaks appear below $r=2d_{s}$. The first peak is at $r=1d_{s}$, which indicates small-small particle pairing, the second indicates small-big particle pairing at $r=1.2d_{s}$, and the third is big-big particle pairing at $r=1.4d_{s}$, which is also the diameter of the big particle.
@@
In Fig.4, panel (a), we observe that for the lowest packing fraction data, the VDOS decreases monotonically with frequency and can be accurately described by a Lorentzian line shape, $g(\omega)=\alpha/(\alpha^{2}+\omega^{2})$ (dashed black line), at least for frequencies below 5 Hz. This line shape is indicative of purely Langevin diffusive dynamics. It is typical of a gas-like state, where particle collisions are almost uncorrelated and independent and can be described by kinetic theory. The Lorentzian fit becomes less accurate at higher frequencies, indicating that the low-packing fraction system is a dilute liquid rather than an ideal gas of free particles. Additionally, the longitudinal and transverse components of $g(\omega)$ are identical, confirming the emergent isotropy of the low-packing fraction phase. Due to the dilute packing, the constituent particles exhibit random, uncorrelated motion in both amplitude and direction. Individual particles’ motion is collisional, resulting in substantial displacements away from their initial positions.
@@
Moving further to $\phi=0.811$ (panel (c) in Fig.4) results in the disappearance of the Lorentzian gas-like contribution to the VDOS at low frequency. The total VDOS increases monotonically with frequency up to a peak located around $\omega=15$ Hz, which corresponds approximately to the average pseudo-Van Hove energy of the emergent collective longitudinal and transverse excitations, as shown in Fig.5. The zero-frequency values of the total VDOS $g(0)$ exhibit a substantial decrease with an increase in the packing fraction $\phi$. This value correlates with the self-diffusion constant $D$, and a direct relation between $D$ and $g(0)$ can be derived for pure Langevin diffusion. This is consistent with the experimental data for the MSD presented in panel (c) of Fig.2. $g(0)$ is largely dominated by the transverse component, and $g_{L}(0)$ is almost zero at $\phi=0.811$. Additionally, the system becomes strongly anisotropic as the longitudinal and transverse VDOS are significantly different, and higher-energy modes appear in the spectrum up to a frequency of approximately $35$ Hz. The maximum frequency mentioned is determined by the highest acquisition rate of the cameras we use to track the position of particles in our granular matter system, which is 40 Hz. This maximum frequency has no intrinsic physical significance and is unrelated to the driving frequency.
@@
As the packing fraction of the system increases to $\phi=0.822$, the value of $g(0)$ becomes extremely small but still finite, indicating that the system is close to the solid phase but not quite there yet. This is later confirmed by the dispersion of shear waves. The longitudinal and transverse VDOS are linear in frequency up to about 8 Hz, as dictated by Debye’s law in 2D, and also commonly found in bulk liquids [73, 74]. However, a sharper peak appears around 12 Hz, which is attributed to the flattening of the dispersion of the collective modes obtained from the dynamical structure factor (panels (d) and (h) in Fig.5). As a result, particle displacements become small, and granular particles move very little away from their initial positions. The dynamics and corresponding VDOS increasingly resemble those of a dense viscous liquid with enhanced solid-like elastic vibrations or a liquid near melting.